OpenStudy (laddiusmaximus):

@zepdrix if f(x)=(x)sqrt(5-x) find f'x find equations to the curve y=(x)sqrt(5-x) at the points (1,2) and (4,4)

zepdrix (zepdrix):

2 separate tangent line equations? :D hmm ok. If we're able to find the derivative effectively, then this problem will be a piece of cake.

OpenStudy (laddiusmaximus):

are we using chain rule here?

zepdrix (zepdrix):

\[\huge f(x)=x \sqrt{5-x}\] Hmm we have the product of two functions of x (two thingies with x in them). So we'll need to apply the product rule.

zepdrix (zepdrix):

If you don't know the derivative of the square root, then make sure you rewrite it as a fractional exponent before differentiating.

OpenStudy (laddiusmaximus):

question if it was just \[\sqrt{5-x}\] then we would use chain rule right?

zepdrix (zepdrix):

Yes.

zepdrix (zepdrix):

Ooo I finally got level 75, yay! :D lol

OpenStudy (laddiusmaximus):

so we get with the product rule, x(1/2(5-x)^-1/2(-1))+(sqrt(5-x)(1)

OpenStudy (laddiusmaximus):

congrats!

zepdrix (zepdrix):

Ok derivative looks good.

zepdrix (zepdrix):

Now that you've taken the derivative, you might want to put it in a form that's easier to work with before we plug the numbers in. It doesn't matter if you do or not, but I always find square roots easier to read the 1/2's.

OpenStudy (laddiusmaximus):

x(-1/2(5-x)^-1/2)+(sqrt(5-x)

zepdrix (zepdrix):

read THAN 1/2's* typo

OpenStudy (laddiusmaximus):

right so that is x(-1/2(1/sqrt(5-x))+(sqrt(5-x))

OpenStudy (laddiusmaximus):

x(-1/2sqrt{5-x}\] +(sqrt(5-x))

zepdrix (zepdrix):

Equation tool giving you trouble? :) hehe

zepdrix (zepdrix):

So now we just plug in our values yes? c:

OpenStudy (laddiusmaximus):

okay so we plug in the y value for (1,2) and (4,4)?

zepdrix (zepdrix):

We don't care about y values. (remember in the previous problem we did? We used the y value for our final setup. But we don't care about it for any of the rest of the problem). Just plug in the x value :)

zepdrix (zepdrix):

Now if we get to a problem that requires us to differentiate IMPLICITLY, then yes, we'll have to use the y value. But we haven't run into a problem like that yet c:

OpenStudy (laddiusmaximus):

okay so 1/2sqrt(2)+sqrt(2)

zepdrix (zepdrix):

water you doing? D: How did you get a 2 under the square root. Neither 5-1 nor 5-4 produce 2.. hmm

OpenStudy (laddiusmaximus):

okay I had 1(1/2sqrt(5-1)) + (sqrt(5-1)

zepdrix (zepdrix):

oh oh i see, so you didn't have the square root anymore :D k fair enough

OpenStudy (laddiusmaximus):

5-1 is 4 sqrt of 4 is 2 right?

zepdrix (zepdrix):

what happened to the negative sign on the first term? Watch those little things mr/ms! >:O

OpenStudy (laddiusmaximus):

ohhhhhhhhh crap

OpenStudy (laddiusmaximus):

-1/4+ 2

OpenStudy (laddiusmaximus):

9/4

zepdrix (zepdrix):

mmmmm close :\ watch those fractions now! :OO

OpenStudy (laddiusmaximus):

7/4

zepdrix (zepdrix):

yay, so whats the equation for the first tangent line? :D

OpenStudy (laddiusmaximus):

y-2=7/4(x-1)

zepdrix (zepdrix):

yay

OpenStudy (laddiusmaximus):

and then we just plug in 4 for the second tangent line

OpenStudy (laddiusmaximus):

I got three for the slope of the second tangent line which gives us y-4=3(x-4)

zepdrix (zepdrix):

Hmm did you forget the minus sign again? I think the slope is suppose to be -1.

OpenStudy (laddiusmaximus):

yup you are right